# One Communication Round is All It Needs for Federated Fine-Tuning Foundation Models

### v1

@@ 3 Theoretical Analysis of One-Shot Federated Fine-Tuning
Foundation Models are Extremely Smooth ($L_{FM}\ll 1$). In Equation 9, the factor $L$ represents the smoothness of the model, with smaller $L$ implying a smoother model. We argue that pre-trained large FMs are much smoother than small models and thus have much smaller $L$ values. Large FMs are pre-trained on large-scale datasets to obtain general capabilities. During this pre-training process, the parameters of FMs are optimized from the ridges to the basins in the loss landscape. Additionally, as observed in a previous work Ainsworth et al. (2022), wider models have more flattened basins in the loss landscapes. With these pieces of prior knowledge, we hold the contention that the loss landscape in large FM fine-tuning is much flatter and smoother than that in training small models from scratch, resulting in much smaller $L$ values. To verify this argument, we estimate $L$ by $L=\frac{\|\nabla F_{i}(\bm{w}_{x})-\nabla F_{i}(\bm{w}_{y})\|}{\|\bm{w}_{x}-\bm{w}_{y}\|}$. We randomly sample a mini-batch of data in the training datasets and compute the gradient on $\bm{w}^{(0,0)}$ and $\bm{w}^{(T,k)}$ to get $\nabla F_{i}(\bm{w}^{(0,0)})$ and $\nabla F_{i}(\bm{w}^{(T,k)})$. Then we visualize the value of $\frac{\|\nabla F_{i}(\bm{w}^{(0,0)})-\nabla F_{i}(\bm{w}^{(T,k)})\|}{\|\bm{w}^{(0,0)}-\bm{w}^{(T,k)}\|}$ in Figure LABEL:fig:L. According to Figure LABEL:fig:L, FMs (i.e., models to the right of the red dash line) have much smaller $L$ values than small models, which is consistent with our conjecture.
@@
Foundation Models Have Much Smaller Model Updates in Fine-Tuning ($\tau_{FM}\ll 1$). Another crucial distinction in our analysis lies in the different tasks in FL: fine-tuning and training from scratch. Since the fine-tuning task updates the model parameters to adapt to downstream tasks without compromising its performance on the general task, it only slightly updates the model parameters. Therefore, the model parameter updates in the fine-tuning process are much smaller than the pre-trained model parameters, i.e., $\|\bm{w}^{(t,j)}-\bm{w}^{(0,0)}\|\ll\|\bm{w}^{(0,0)}\|$. In this case, the federated fine-tuning task would have a very small $\tau$ in Equation 9. To verify this, we conduct experiments to estimate the $\tau$ values by $\frac{\|\bm{w}^{(T,k)}-\bm{w}^{(0,0)}\|}{\|\bm{w}^{(0,0)}\|}$, where $\bm{w}^{(T,k)}$ represents the model update after the entire fine-tuning process on the training datasets. We visualize the estimated $\tau$ values of different models in Figure LABEL:fig:tau, which illustrates that the $\tau$ values in FMs are much smaller than those in small models.
@@
Large Foundation Models Require Less Fine-Tuning Epochs ($Tk_{FM}\ll Tk_{small}$). Different from training a small model from scratch, fine-tuning a large model typically doesn’t require a large number of total training steps to ensure convergence. This is mainly because the pre-trained models will be overfitting on the fine-tuning data with too many epochs, which will destroy the model’s ability on the general tasks. As a result, the $Tk$ values of large FMs are also smaller than those in small models. Table. 4 in the Appendix displays the $T$ and $k$ numbers adopted by our experiments.
@@
We also visualize $\|\bm{w}^{(0,0)}\|$ in Figure LABEL:fig:w00. Although the $\|\bm{w}^{(0,0)}\|$ value of the small model is relatively small, it does not exhibit a clear trend positively correlated with model size (e.g., TinyLlama has a similar $\|\bm{w}^{(0,0)}\|$ value with BERT, but has 10 times more parameters than BERT, Gemma-2b has much larger $\|\bm{w}^{(0,0)}\|$ value than Llama-13b).
@@
Large Foundation Models Have Smaller One-Shot Federated Fine-tuning Error $\varepsilon$. Based on the discussion before regarding the $L$, $\tau$, $Tk$, and $\|\bm{w}^{(0,0)}\|$ values of the model with various sizes, we conclude that large FMs have smaller $L$, $\tau$, and $Tk$ values, while $\|\bm{w}^{(0,0)}\|$ is not strongly related to the model size. We finally visualize the $\|\varepsilon\|=\Gamma\|\bm{w}^{(0,0)}\|$ values of different models in Figure 4. The results in Figure 4 clearly demonstrate that large FMs (GPT-2 and all models to its right) have significantly lower $\|\varepsilon\|$ values than the small models, with larger FMs having lower values. According to Eq. 9, smaller $\|\varepsilon\|$ means a smaller difference between one-shot and multi-round FL. Consequently, FMs have much better one-shot FL performance than small models. The larger FM has lower errors in one-shot federated fine-tuning.
@@
In summary, there are three main reasons why FMs have smaller errors in one-shot federated fine-tuning. First, the pre-trained FMs have extremely smooth loss landscapes in fine-tuning, i.e., $L_{FM}\ll 1$. Second, the fine-tuning model updates are particularly small compared to the pre-trained parameters, i.e., $\tau_{FM}\ll 1$. Third, FM fine-tuning requires far fewer epochs than training small models from scratch, i.e., $Tk_{FM}\ll Tk_{small}$. These three factors lead to much smaller error $\varepsilon$ in one-shot federated fine-tuning of FMs.
4 Experiment
4.1 Experimental Setups
@@
Models and Datasets. To demonstrate the performance of FMs in different sizes, we selected multiple models ranging in parameter size from 1b to 13b for experiments. The language FMs we experimented with range in parameter size from smallest to largest as follows: TinyLlama (1.1b) Zhang et al. (2024b), Gemma-2b Team et al. (2024), Llama-7b, and Llama-13b Touvron et al. (2023a). We use the MMLU Hendrycks et al. (2020) training dataset and Wizard Luo et al. (2023) dataset to federated fine-tune these models. For evaluation, we leverage MMLU and ARC Challenge Clark et al. (2018) in Eval-Harness Gao et al. (2023) to evaluate the model ability of QA tasks, and the GPT-4 evaluation in MT-bench Zheng et al. (2023) for the chat assistant task.
Federated Fine-Tuning Settings. For federated fine-tuning on a single MMLU or Wizard dataset, we randomly split the dataset into 10 clients. We also have a strongly non-iid setting, which assigns the MMLU dataset to 10 clients and the Wizard dataset to another 10 clients, and lets the 20 clients fine-tune the FM. For the baseline, we use a multi-round FedAvg algorithm on both LoRA and full fine-tuning. For our one-shot federated fine-tuning, we only perform a single communication round. To ensure fairness, we keep the total number of local epochs the same between multi-round and one-shot federated fine-tuning. e.g., , if the setting in multi-round federated fine-tuning is 3 communication rounds, 1 local epoch in each round, the setting in one-shot should be 1 communication round, 3 local epoch in that round.
4.2 Main Results
One-Shot Federated Fine-Tuning in QA Tasks. We first evaluate the performance of one-shot federated fine-tuning in QA tasks and display the results in Table 1. The columns with titles MMLU, Wizard, and M-W represent the model fine-tuned by MMLU, Wizard, and the mixture of MMLU and Wizard datasets respectively. The rows with the title MMLU and ARC represent the model accuracy evaluated by the MMLU test set and ARC Challenge. The Methods columns mean the fine-tuning is performed by LoRA or full fine-tuning, while the rows with a star (*) represent one-shot federated fine-tuning. According to Table 1, the performance of one-shot federated fine-tuning is generally comparable to that of multi-round federated fine-tuning. In some settings, one-shot fine-tuning achieves higher accuracy. For example, the Llama-13b one-shot fine-tuned by LoRA on the Wizard dataset achieves 47.93% accuracy on MMLU and 58.11% on ARC Challenge, which is higher than the 46.83% and 55.72% accuracy of multi-round fine-tuning. In full fine-tuning, multi-round fine-tuning performs better in some settings. For instance, the Llama-13b multi-round full fine-tuned on the Wizard dataset outperforms one-shot fine-tuning on both MMLU and ARC Challenge. These observations align with our previous theoretical analysis. Full fine-tuning involves greater parameter updates compared to LoRA, resulting in a larger $\tau$ value, and thus a larger $\varepsilon$ value. Consequently, the performance of one-shot full fine-tuning may sometimes be inferior to LoRA fine-tuning. However, this does not affect our overall conclusion: for FMs, one-shot federated fine-tuning can effectively replace multi-round federated fine-tuning. One-shot fine-tuning provides comparable performance to multi-round fine-tuning while significantly reducing communication costs.
@@ 4.2 Main Results
Table 1: Performance of one-shot federated fine-tuning in Q&A tasks. The rows with star (*) are the results of one-shot federated fine-tuning.
@@
One-Shot Federated Fine-Tuning in Chat Assistant Tasks. We evaluate the performance of FMs in chat assistant tasks, where models generate answers to several questions and are scored by GPT-4. The score from MT-bench is the average score across all questions. Table 2 shows the scores of multi-round and one-shot federated fine-tuned models. The averaging scores of three fine-tuning datasets indicate that larger FMs perform better in one-shot federated fine-tuning. Specifically, multi-round fine-tuning outperforms one-shot fine-tuning in both LoRA and full fine-tuning on the Tinyllama model, which is the smallest model in our experiments. On the contrary, for larger models, such as Gemma-7b and Llama-13b, one-shot fine-tuning performs better than multi-round fine-tuning. This observation aligns with our previous theoretical analysis that larger models have smaller one-shot fine-tuning errors. The superior performance of one-shot fine-tuning in larger models might be attributed to the larger number of local epochs per round, which leads to a slower local learning rate decay. The chat assistant’s capabilities may benefit from this smoother learning rate decay process.
@@
Table 2: Performance of one-shot federated fine-tuning on chat assistant tasks. Wizard has better performance than MMLU on MT-bench. We use AVG. column to show the averaging performance of specific methods.
Figure 5: "A photo of a dog in a bucket" generated by LoRA fine-tuned stable diffusion models.
One-Shot Federated Fine-Tuning in Text-To-Image Generation Tasks. In addition to testing LLMs, we also evaluated the effectiveness of one-shot federated fine-tuning in the text-to-image generation tasks. We use LoRA to fine-tune a stable-diffusion-v1-5 Rombach et al. (2022) on the Dreambooth Ruiz et al. (2023) dataset with 5 distributed clients. In the multi-round setting, we have 5 global rounds, with 5 local epochs in each round. In the one-shot setting, we have 1 global round with 25 local epochs. After fine-tuning, we evaluated the models by the CLIP Hessel et al. (2021) score with ViT-B-32 Dosovitskiy et al. (2020) to assess the quality of generated images. Figure 5 shows the images generated with the prompt "A photo of a dog in a bucket" The right column displays the result of multi-round federated fine-tuning, while the left column shows the result from the one-shot setting. The numbers to the right of the images represent the CLIP scores. The qualities of the images generated by both methods are essentially the same. The average CLIP score in the one-shot setting is 0.3343, while the score in the multi-round setting is 0.3341. These results indicate that the effectiveness of one-shot federated fine-tuning extends to fine-tuning stable diffusion models.
5 Discussion
One-Shot Federated Fine-Tuning Saves Communication Cost.
In FL, the server sends the model parameters to all the selected clients and receive the clients’ model updates in each communication round. Thus, the total number of communicated parameters in multi-round should be $2mTS$, where $S$ is the model size. In one-shot federated fine-tuning, the server and the clients only perform one-round communication, so the number of communicated parameters is only $2mS$. This reduction in communication overhead is significant, especially when fine-tuning large FMs. For instance, the Llama-13b model has approximately 50GB parameters, i.e., $S=50GB$. In our experiments, the 3-round federated fine-tuning on Llama-13b needs to communicate 3000GB data between the server and the clients, which may be unaffordable in scenarios with tight communication budgets. However, one-shot federated fine-tuning reduces this amount to 1000GB. This substantial reduction in communication makes federated fine-tuning of large FMs more practical and affordable in real-life scenarios.
One-Shot Federated Fine-Tuning Supports Asynchronous Global Aggregation.
In traditional multi-round FL, clients need to train local models synchronously. The server can only perform the aggregation and send the new global model to clients after receiving all local model updates. This requirement poses challenges for federated learning applications. For example, if local computation resources are occupied by other tasks or if the connection between the server and clients is unstable, the training process will be halted. One-shot federated fine-tuning effectively addresses this problem. The server can update the global model with local updates as soon as they are received, allowing for real-time model updates. Therefore, even if some clients fail to send model updates promptly due to various reasons, the global model on the server can still be updated by most clients, resulting in a usable global model. To further illustrate this point, we sequentially aggregated local model updates from client 1 to client 10 in one-shot federated fine-tuning of Llama-7b on the Wizard dataset. We tested the global model’s performance on the MT-bench as we aggregated updates from 1, 2, 3, …, and up to 10 clients. The results are displayed in Figure 6. The model score increases as more clients contribute their local updates to the global model, indicating that each individual local model update provides an immediate improvement in global model performance. The red dash line represents the model score in the synchronous FL setting, which is equal to the score of aggregating ten clients in asynchronous FL.
One-Shot Federated Fine-Tuning Naturally Mitigates Client-Side Privacy Threatens.
In traditional FL algorithms, clients repeatedly receive new global model parameters each round, which could lead to client-side privacy issues. Malicious clients can exploit model inversion Fredrikson et al. (2015); Zhang et al. (2020) and gradient inversion attacks Huang et al. (2021) to recover private training samples or user inputs from other clients Wei et al. (2023). These attacks heavily rely on access to the global model parameters and certain data distribution information. However, in one-shot FL, the server can choose not to send back global parameters and only provide an API of the fine-tuned model. By doing this, it can eliminate the possibility of client-side privacy leakage.
6 Conclusion
In this paper, we tackle the critical issue of high communication costs that limit the practical application of federated fine-tuning. Through a series of experiments, we demonstrate that multi-round communication is not necessary for fine-tuning FMs, as one-shot federated fine-tuning achieves comparable performance. We then provide a theoretical analysis to explain why one-shot federated fine-tuning is effective for large FMs and validate our findings with empirical evidence. Our extensive experiments show that one-shot federated fine-tuning performs on par with multi-round federated fine-tuning across 5 different models and 3 diverse tasks. This method significantly reduces communication overhead, making federated fine-tuning more feasible and efficient, especially for large-scale models. Moreover, one-shot federated fine-tuning supports asynchronous local updates and enhances security by minimizing data exposure during the training process. These findings make it possible to harness the power of large FMs in environments with limited communication resources, thereby broadening the accessibility and utility of advanced AI technologies.
Appendix A Related Work
One-Shot Federated Learning.
One-shot federated learning refers to learning the parameters of the global model in a single round of communication between clients and the server (Guha et al., 2019). There are two main strategies for optimizing one-shot FL, neuron matching and knowledge distillation. Neuron matching is based on the permutation symmetry of neural networks (Ainsworth et al., 2022), which means that client model parameters can be aligned according to a common ordering and then be averaged. Previous works use algorithms such as the Fisher information matrix (Jhunjhunwala et al., 2024) and permutation matrix (Wang et al., 2020a) to match the local model parameters. The knowledge distillation methods aim at distilling knowledge from well-trained local models through public data (Gong et al., 2021; Li et al., 2020; Heinbaugh et al., 2022). Some works also use distilled data to transfer knowledge between clients and the server (Zhou et al., 2020). Recent works adopt generative models to help generate substitute data for the local dataset on the server (Yang et al., 2024; Zhang et al., 2022).
Federated Fine-Tuning.
Federated fine-tuning (Orescanin et al., 2021; Cheng et al., 2021) aims to fine-tune FMs by cross-domain on-device datasets while preserving data privacy. Recent works use PEFT methods such as LoRA (Hu et al., 2021) in federated fine-tuning (Zhang et al., 2024a) to save communication and computation costs. Federated fine-tuning also faces similar research problems as FL. Current works have discussed the non-IID problem (Cho et al., 2024) and personalized federated fine-tuning (Wagner et al., 2024).
Appendix B Additional Experimental Setups
Computer Resources.
We used a 256GB AMD EPYC 7763 64-Core Processor on Linux v4.18.0 to run the experiments. For LoRA fine-tuning on all the models and full fine-tuning on all the models except Llama-13b, we used 4 NVIDIA RTX A6000 GPUs. For Llama-13b full fine-tuning, we use 8 NVIDIA A100 GPUs.
Hyperparameter Settings.
@@ Appendix B Additional Experimental Setups
For LoRA fine-tuning across all the models and datasets, we set the local LoRA rank to 16, the local learning rate to 3e-4, and the batch size to 64. For full fine-tuning, we reduced the learning rate to 3e-5 and set the learning rate to 8. For multi-round settings, the numbers of global communication rounds and local epochs in each round in different models and datasets are listed in Table 3. The one-shot setting satisfies $T=1$ and $k$ equals $Tk$ in the multi-round setting. The number of rounds and epochs we selected can ensure convergence and avoid overfitting. We show a simple example in Appendix C to demonstrate this point.
@@
Table 3: Global rounds and local epochs settings in multi-round experiments.
@@
Appendix C Additional Experimental Results
Table 4: $Tk$ settings in experiments. $T$ is the number of global communication rounds. $k$ is the total number of local SGD steps, which is computed by (dataset length $\times$ epoch number /// batch size).
T 50 50 5 3 3 3 3 3
k 7812 3906 5625 3750 1875 1875 1875 1875
Tk 390600 195300 28125 11250 5625 5625 5625 5625
Zero-Shot Results.
We test the zero-shot performance of models used in Table 1 for reference. The results are displayed in Table 5
Table 5: Zero-Shot results of models on MMLU and ARC Challenge.

### v2

@@ III Theoretical Analysis of One-Shot Federated Fine-Tuning
Foundation Models are Smoother than Small Models ($L_{FM}\ll L_{SM}$). In Equation 9, the factor $L$ represents the smoothness of the model, with smaller $L$ implying a smoother model. We argue that pre-trained large FMs are much smoother than small models and thus have much smaller $L$ values. FMs are pre-trained on large-scale datasets to obtain general capabilities. During this pre-training process, the parameters of FMs are optimized from the ridges to the basins in the loss landscape. Additionally, as observed in a previous work [ainsworth2022git], wider models have more flattened basins in the loss landscapes. With these pieces of prior knowledge, we hold the contention that the loss landscape in large FM fine-tuning is much flatter and smoother than that in training small models from scratch, resulting in much smaller $L$ values. To verify this argument, we estimate $L$ by $L=\frac{\|\nabla F_{i}(\bm{w}_{x})-\nabla F_{i}(\bm{w}_{y})\|}{\|\bm{w}_{x}-\bm{w}_{y}\|}$. We randomly sample a mini-batch of data in the training datasets and compute the gradient on $\bm{w}^{(0,0)}$ and $\bm{w}^{(T,k)}$ to get $\nabla F_{i}(\bm{w}^{(0,0)})$ and $\nabla F_{i}(\bm{w}^{(T,k)})$. Then we visualize the value of $\frac{\|\nabla F_{i}(\bm{w}^{(0,0)})-\nabla F_{i}(\bm{w}^{(T,k)})\|}{\|\bm{w}^{(0,0)}-\bm{w}^{(T,k)}\|}$ in Fig. 2a. According to Fig. 2a, FMs (i.e., models to the right of the red dash line) have much smaller $L$ values than small models, which is consistent with our conjecture.
@@
Foundation Models Have Much Smaller Model Updates in Fine-Tuning ($\tau_{FM}\ll\tau_{SM}$). Another crucial distinction in our analysis lies in the different tasks in FL: fine-tuning and training from scratch. Since the fine-tuning task updates the model parameters to adapt to downstream tasks without compromising its performance on the general task, it will only slightly update the model parameters. Therefore, the model parameter updates in the fine-tuning process are much smaller than the pre-trained model parameters, i.e., $\|\bm{w}^{(t,j)}-\bm{w}^{(0,0)}\|\ll\|\bm{w}^{(0,0)}\|$. In this case, the federated fine-tuning task would have a very small $\tau$ in Equation 9. To verify this, we conduct experiments to estimate the $\tau$ values by $\frac{\|\bm{w}^{(T,k)}-\bm{w}^{(0,0)}\|}{\|\bm{w}^{(0,0)}\|}$, where $\bm{w}^{(T,k)}$ represents the model update after the entire fine-tuning process on the training datasets. We visualize the estimated $\tau$ values of different models in Fig. 2b, which illustrates that the $\tau$ values in FMs are much smaller than those in small models.
(a) One-Shot Aggregation
(b) Multi-Round Aggregation of FMs
(c) Multi-Round Aggregation of SMs
Figure 3: Aggregation in FL with two clients. (a): one round aggregation with $Tk$ local epochs of both the FM and the small model (SM). Although the FM and the SM may differ in practice, we depict both using the same setup as in (a) to make the comparison easier to understand. (b): FL of FM under $T$ round aggregation with $k$ local epochs, where the landscape is smooth and fine-tuning makes small changes, so the new local directions remain close to the one-shot direction. (c): FL of SM under $T$ round aggregation with $k$ local epochs, where the training direction changes markedly after each global update. The final global model deviates significantly from the one-shot global model. This figure visualizes the intuition behind our analysis that one round works well for FMs.
@@
Large Foundation Models Require Less Fine-Tuning Steps ($Tk_{FM}\ll Tk_{small}$). Different from training a small model from scratch, fine-tuning a large model typically doesn’t require a large number of total training steps to ensure convergence. This is mainly because the pre-trained models will be overfitting on the fine-tuning data with too many epochs, which will destroy the model’s ability on the general tasks. As a result, the $Tk$ values of large FMs are also smaller than those in small models. Table. I displays the $T$ and $k$ numbers adopted by our experiments.
@@
We also visualize $\|\bm{w}^{(0,0)}\|$ in Fig. 2c. Although the $\|\bm{w}^{(0,0)}\|$ value of the small model is relatively small, it does not exhibit a clear trend positively correlated with model size (e.g., TinyLlama has a similar $\|\bm{w}^{(0,0)}\|$ value with BERT, but has 10 times more parameters than BERT, Gemma-2b has much larger $\|\bm{w}^{(0,0)}\|$ value than Llama-13b).
@@
Conclusion: Large Foundation Models Have Smaller Global Difference $\varepsilon$. Based on the discussion before regarding the $L$, $\tau$, $Tk$, and $\|\bm{w}^{(0,0)}\|$ values of the model with various sizes, we conclude that large FMs have smaller $L$, $\tau$, and $Tk$ values, while $\|\bm{w}^{(0,0)}\|$ is not strongly related to the model size. We ignore the client number $m$ and visualize the $\|\varepsilon\|=\Gamma\|\bm{w}^{(0,0)}\|$ values of different models in Fig. 4. The results in Fig. 4 clearly demonstrate that large FMs (GPT-2 and all models to its right) have significantly lower $\|\varepsilon\|$ values than the small models, with larger FMs having lower values. According to Eq. 9, smaller $\|\varepsilon\|$ means a smaller difference between one-shot and multi-round FL. Consequently, FMs have much better one-shot FL performance than small models.
@@
In summary, the reasons why large FMs have smaller differences in one-shot federated fine-tuning are due to three main factors. First, the pre-trained FMs have extremely smooth loss landscapes in fine-tuning, i.e., $L_{FM}\ll L_{SM}$. Second, the fine-tuning model updates are particularly small compared to the pre-trained parameters, i.e., $\tau_{FM}\ll\tau_{SM}$. Third, FM fine-tuning requires far fewer epochs than training small models from scratch, i.e., $Tk_{FM}\ll Tk_{SM}$. These three factors lead to much smaller error $\varepsilon$ in the one-shot federated fine-tuning of FMs.
Our theoretical analysis focuses on the error gap between one-shot and standard multi-round FL, namely the discrepancy caused by using a single communication round, rather than on asymptotic convergence. Unlike classical FL results that prove convergence as $T\!\to\!\infty$ [wang2020tackling, kairouz2021advances], Theorem 1 expresses this gap as an explicit function of model- and task-dependent quantities (e.g., smoothness, update magnitude, and training horizon), thereby predicting how model scale and task regime (fine-tuning v.s. training from scratch) influence one-shot FL. These predictions are borne out in our experiments, whereas asymptotic analyses do not capture such finite-round, size- and task-sensitive effects.
These proofs and experiments also support the following intuition: because the loss landscape of large pre-trained models is relatively smooth and fine-tuning induces only small parameter changes, the update directions remain largely stable even when the clients’ starting points shift across rounds in multi-round FL. As a result, when we switch to one-shot FL and do not refresh the starting point, the effective direction remains close to that of multi-round FL, which yields an outcome similar to vector aggregation. In contrast, smaller models encounter a more rugged loss landscape, so clients require multiple FL rounds to repeatedly adjust the update direction and search for a better optimum. Under one-shot FL, each client continues to fine-tune from the initial parameters throughout, which causes error accumulation and ultimately degrades performance. Fig. 3 describes this intuition. The error of one-shot FL in small model accumulates with the multi-round global aggregation (Fig. 3 (c)), while the error of FMs (Fig. 3 (b)) remains small due to the smooth loss landscape and small fine-tuning updates.
IV Experiment
IV-A Experimental Setups
@@
Models and Datasets. To demonstrate the performance of FMs of different sizes on one-shot federated fine-tuning, we selected multiple models ranging in parameter size from 1b to 13b for experiments. The language models we experimented with range in parameter size from smallest to largest as follows: TinyLlama (1.1b) [zhang2024tinyllama], Gemma-2b, Gemma-7b [team2024gemma], Llama-7b, and Llama-13b [touvron2023llama]. We use the MMLU [hendrycks2020measuring] training dataset and Wizard [luo2023wizardmath] dataset to federated fine-tune these models. For evaluation, we leverage MMLU and ARC Challenge [allenai:arc] in Eval-Harness [eval-harness] to evaluate the model ability of QA tasks, and the GPT-4 evaluation in MT-bench [zheng2023judging] for the chat assistant task.
TABLE I: TkTk settings in experiments. $T$ is the number of global communication rounds. $k$ is the total number of local SGD steps, which is computed by (dataset length $\times$ epoch number // batch size).
T 50 50 5 3 3 3 3 3
k 7812 3906 5625 3750 1875 1875 1875 1875
Tk 390600 195300 28125 11250 5625 5625 5625 5625
@@
TABLE II: Performance of multi-round and one-shot federated fine-tuning in Q&A tasks. The rows with star (*) are the results of one-shot federated fine-tuning.
@@
TABLE III: Performance of one-shot federated fine-tuning on chat assistant tasks. Wizard has better performance than MMLU on MT-bench. We use AVG. column to show the averaging performance of specific methods.
@@
Federated Fine-Tuning Settings. For federated fine-tuning on a single MMLU or Wizard dataset, we randomly split the dataset into 10 clients. We also have a strongly non-iid setting, which assigns the MMLU dataset to 10 clients and the Wizard dataset to another 10 clients, and lets the 20 clients fine-tune the FM. For the baseline, we use a multi-round FedAvg algorithm on both LoRA and full fine-tuning. For our one-shot federated fine-tuning, we only perform a single communication round. To ensure fairness, we keep the total number of local steps the same between multi-round and one-shot federated fine-tuning. e.g., , if the setting in multi-round federated fine-tuning is 3 communication rounds, 1 local epoch in each round, the setting in one-shot should be 1 communication round, 3 local epochs in that round. For LoRA fine-tuning across all the models and datasets, we set the local LoRA rank to 16, the local learning rate to 3e-4, and the batch size to 64. For full fine-tuning, we reduced the learning rate to 3e-5 and set the batch size to 8. For multi-round settings, the numbers of global communication rounds and local epochs in each round in different models and datasets are listed in Table IV. The one-shot setting satisfies $T=1$ and $k$ equals $Tk$ in the multi-round setting. The number of rounds and epochs we selected can ensure convergence and avoid overfitting. We show a simple example in Fig. 7 to demonstrate this point. Note that, given the inherent heterogeneity of text datasets (e.g., differing sequence lengths), heterogeneity is present in all our experiments.
Computer Resources.For LoRA fine-tuning on all the models and full fine-tuning on all the models except Llama-13b, we used 4 NVIDIA RTX A6000 GPUs. For Llama-13b full fine-tuning, we use 8 NVIDIA A100 GPUs.
@@
TABLE IV: Global rounds and local epochs settings in multi-round experiments.
@@
IV-B Main Results
One-Shot Federated Fine-Tuning in QA Tasks. We first evaluate the performance of one-shot federated fine-tuning in QA tasks and display the results in Table II. The columns with titles MMLU, Wizard, and M-W represent the model fine-tuned by MMLU, Wizard, and the mixture of MMLU and Wizard datasets respectively. The rows with the title MMLU and ARC represent the model accuracy evaluated by the MMLU test set and ARC Challenge. The Methods columns mean the fine-tuning is performed by LoRA or full fine-tuning, while the rows with a star (*) represent one-shot federated fine-tuning. According to Table II, the performance of one-shot federated fine-tuning is generally comparable to that of multi-round federated fine-tuning. In some settings, one-shot fine-tuning even achieves higher accuracy. For example, the Llama-13b model one-shot fine-tuned by LoRA on the Wizard dataset achieves 47.93% accuracy on MMLU and 58.11% on ARC Challenge, which is higher than the 46.83% and 55.72% accuracy of multi-round fine-tuning. In full fine-tuning, multi-round fine-tuning performs better in some settings. For instance, the Llama-13b model multi-round full fine-tuned on the Wizard dataset outperforms one-shot fine-tuning on both MMLU and ARC Challenge. These observations align with our previous theoretical analysis. Full fine-tuning involves greater parameter updates compared to LoRA, resulting in a larger $\tau$ value, and thus a larger $\varepsilon$ value. Consequently, the performance of one-shot full fine-tuning may sometimes be inferior to LoRA fine-tuning. However, this does not affect our overall conclusion: for large FMs, one-shot federated fine-tuning can effectively replace multi-round federated fine-tuning. One-shot fine-tuning provides comparable performance to multi-round fine-tuning while significantly reducing communication costs. To clarify the fine-tuning gain, we test the zero-shot performance of models used in Table II for reference. The results are displayed in Table V.
One-Shot Federated Fine-Tuning in Chat Assistant Tasks. We evaluate the performance of FMs in chat assistant tasks, where models generate answers to several questions and are scored by GPT-4. The score from MT-bench is the average score across all questions. Table III shows the scores of multi-round and one-shot federated fine-tuned models. The averaging scores of three fine-tuning datasets indicate that larger FMs perform better in one-shot federated fine-tuning. Specifically, multi-round fine-tuning outperforms one-shot fine-tuning in both LoRA and full fine-tuning on the TinyLlama model, which is the smallest model in our experiments. On the contrary, for larger models, such as Gemma-7b and Llama-13b, one-shot fine-tuning performs better than multi-round fine-tuning. This observation aligns with our previous theoretical analysis that larger models have smaller differences between one-shot and multi-round FL. The superior performance of one-shot fine-tuning in larger models might be attributed to the less number of local epochs per round, which leads to a slower local learning rate decay. The chat assistant’s capabilities may benefit from this smoother learning rate decay process.
Figure 5: “A photo of a dog in a bucket” generated by LoRA fine-tuned stable diffusion models. The left column shows images generated by the one-shot FL models, along with their CLIP scores; the right column shows the results of multi-round FL.
One-Shot Federated Fine-Tuning in Text-To-Image Generation Tasks. In addition to testing LLMs, we also evaluated the effectiveness of one-shot federated fine-tuning in the text-to-image generation tasks. We use LoRA to fine-tune a stable-diffusion-v1-5 [rombach2022high] model on the Dreambooth [ruiz2023dreambooth] dataset with 5 distributed clients. In the multi-round setting, we have 5 global rounds, with 5 local epochs in each round. In the one-shot setting, we have 1 global round and 25 local epochs in that round. After fine-tuning, we evaluated the models using the CLIP [hessel2021clipscore] score with ViT-B-32 [dosovitskiy2020image] to assess the quality of generated images based on specific prompts. Fig. 5 shows the images generated with the prompt ”A photo of a dog in a bucket” The right column displays the result of multi-round federated fine-tuning, while the left column shows the result from the one-shot setting. The numbers to the right of the images represent the CLIP scores. The qualities of the images generated by both methods are essentially the same. The average CLIP score in the one-shot setting is 0.3343, while the score in the multi-round setting is 0.3345. These results indicate that the effectiveness of one-shot federated fine-tuning extends to fine-tuning stable diffusion models.
TABLE V: Zero-Shot results of models on MMLU and ARC Challenge.